OpenStudy (nonopro):

Lucas cut a board that was 3 1/5 in. long and 3 1/3 in. wide. What was the area of the board? Express your answer in simplest form. A.15 1/5 in² B.10 2/3 in² C.9 2/5 in² D.6 8/15 in²

OpenStudy (anonymous):

\(\color {olive} {\text{First know the formula of Area of a Rectangle:}}\) \(A=length \times width\) \(\color {olive} {\text{Change the the given (mixed numbers) to improper fractions:}}\) \(l=3\frac{1}{5} \rightarrow \frac{16}{5}\) \(w=3\frac{1}{3} \rightarrow \frac{10}{3}\) \(\color {olive} {\text{Plug in the formula the given:}}\) \(\color {indigo} {\text{GIVEN:}}\) \(l=\frac{16}{5}\) \(w=\frac{10}{3}\) \(\huge A=\frac{16}{5} \times \frac{10}{3}\) \(\color {olive} {\text{Multiply the numerators and denominators:}}\) \(\huge A=\frac{16}{5} \times \frac{10}{3} =\frac{16 \times 10}{5 \times 3}=\frac {160}{15}\)

OpenStudy (jdoe0001):

well, is a board, so I assume a rectangle thus area of a rectangle = length * width multiply the given measures

OpenStudy (anonymous):

\(\color {olive} {\text{Now simplify}\huge \frac{160}{15}}.\)